OpenStudy (anonymous):

PLEASE HELP!!!!!! Sam is observing the velocity of a car at different times. After two hours, the velocity of the car is 50 km/h. After six hours, the velocity of the car is 54 km/h. Part A: Write an equation in two variables in the standard form that can be used to describe the velocity of the car at different times. Show your work and define the variables used. Part B: How can you graph the equation obtained in Part A for the first seven hours?

OpenStudy (anonymous):

@nikato @jim_thompson5910 @mathslover

jimthompson5910 (jim_thompson5910):

are you able to get started?

OpenStudy (anonymous):

ya

jimthompson5910 (jim_thompson5910):

what do you have so far

OpenStudy (anonymous):

i don't know what to do

jimthompson5910 (jim_thompson5910):

but you said you were able to get started?

OpenStudy (anonymous):

i thought you meant was i able 2 get started on the problem with u

jimthompson5910 (jim_thompson5910):

ok let's start off defining the variables

jimthompson5910 (jim_thompson5910):

what are the two variables?

OpenStudy (anonymous):

idk

jimthompson5910 (jim_thompson5910):

read the problem again and think it over what two things are we trying to connect?

OpenStudy (anonymous):

the velocity of 2 and 6 hours?

jimthompson5910 (jim_thompson5910):

velocity is one variable what's the other?

OpenStudy (anonymous):

the car?

jimthompson5910 (jim_thompson5910):

no, what's another thing that has some numerical values to it

OpenStudy (anonymous):

the time?

jimthompson5910 (jim_thompson5910):

good, the two variables are time and speed (velocity)

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

so let's define x = time y = speed

jimthompson5910 (jim_thompson5910):

"After two hours, the velocity of the car is 50 km/h" that means x = 2 y = 50

jimthompson5910 (jim_thompson5910):

giving us the point (2,50)

jimthompson5910 (jim_thompson5910):

what's the other point?

OpenStudy (anonymous):

(6,54)?

jimthompson5910 (jim_thompson5910):

now find the slope of the line through the two points (2,50) and (6,54)

OpenStudy (anonymous):

how do i do that?

jimthompson5910 (jim_thompson5910):

using the slope formula m = (y2 - y1)/(x2 - x1)

jimthompson5910 (jim_thompson5910):

basically, you subtract the y coordinates then subtract the x coordinates (same order) finally, divide to get the slope (y over x)

OpenStudy (anonymous):

I got 1

jimthompson5910 (jim_thompson5910):

me too

jimthompson5910 (jim_thompson5910):

so m = 1

jimthompson5910 (jim_thompson5910):

Now plug m = 1, x = 2 and y = 50 into y = mx + b and solve for b y = mx + b y = 1*x + b 50 = 1*2 + b ... ... ... b = ???

OpenStudy (anonymous):

48?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so the equation is y = 1x + 48 or just y = x + 48

jimthompson5910 (jim_thompson5910):

notice how when x = 2, y is... y = x + 48 y = 2 + 48 y = 50 so when the time is 2 hrs, the speed is 50 km/hr

jimthompson5910 (jim_thompson5910):

also, notice how when x = 6, y is... y = x + 48 y = 6 + 48 y = 54 so when the time is 6 hrs, the speed is 54 km/hr

OpenStudy (anonymous):

ok, but how do i put that in standard form like the question asks?

jimthompson5910 (jim_thompson5910):

now you must convert y = x + 48 into the form Ax + By = C

OpenStudy (anonymous):

how?

jimthompson5910 (jim_thompson5910):

by getting that x over to the left side

OpenStudy (anonymous):

can u show me how 2 do that?

jimthompson5910 (jim_thompson5910):

think of x+48 as 48+x

jimthompson5910 (jim_thompson5910):

y = 48+x how do we move that +x over? how do we get rid of it on that right side?

OpenStudy (anonymous):

by subtracting it?

jimthompson5910 (jim_thompson5910):

good